Let an denote the number of surjective (onto) functions f : {1, 2, . . . , n} −→

{1, 2, 3} such that f(1) < f(2). Give a Θ estimate for an.


1
Expert's answer
2021-12-30T17:14:51-0500

number of surjective functions "f:[k]\to[n]" :

"N=\displaystyle \sum_{i=0}^n \begin{pmatrix}
 n \\
 i
\end{pmatrix}(-1)^i(n-i)^k"


in our case:

"N=\displaystyle \sum_{i=0}^n \begin{pmatrix}
 3 \\
 i
\end{pmatrix}(-1)^i(3-i)^n=3^n-3\cdot2^n+3"


minimum number of function such that f(1) > f(2) is 0

maximum number of function such that f(1) > f(2) is 3:

"f(2)=1,f(2)=2,f(1)=3"


so, "\theta" estimate for an :

"N-3\le a_n \le N"

"3^n-3\cdot2^n<a_n<3^n-3\cdot2^n+3"